Two 2.30cm×2.30cm plates that form a parallel-plate capacitor are charged to ± 0.708 nC .

Part A

What is the potential difference across the capacitor if the spacing between the plates is 0.800 mm ?

Express your answer with the appropriate units.

Part B

What is the potential difference across the capacitor if the spacing between the plates is 1.60 mm ?

Express your answer with the appropriate units.
Answer #1

Part A.

Charge on capacitor is given by:

Q = C*V

So potential difference across capacitor will be

V = Q/C

Q = Charge = 0.708 nC = 0.708*10^-9 C

C = capacitance = e0*A/d

e0 = 8.85*10^-12

A = Area of plates = 2.30*2.30 cm^2 = 5.29*10^-4 m^2

d = spacing between plates = 0.800 mm = 0.800*10^-3 m

So,

V = Q/C = Q*d/(e0*A)

Using given values:

V = 0.708*10^-9*0.800*10^-3/(8.85*10^-12*5.29*10^-4) = 120.98 V

V = Potential difference = 121 V

Part B.

when d = 1.60 mm = 1.60*10^-3 m

So Now,

V = 0.708*10^-9*1.60*10^-3/(8.85*10^-12*5.29*10^-4) = 241.96 V

V = Potential difference = 242 V

Please Upvote.

Comment below if you've any query.